OpenStudy (anonymous):

2.Solve M = 2x + 3y for y

jhonyy9 (jhonyy9):

M-2x=3y y=(M-2x)/3

OpenStudy (anonymous):

\[M-2x=3y\]\[y=\frac{M-2x}{3}\]

OpenStudy (anonymous):

thank u